Suppose a random variable  x is best described by a normal distribution with  μ = 60 and   Find the  z-score that corresponds to the value  x = 63.


1
Expert's answer
2022-03-29T17:46:24-0400

"Z=\frac{x-\mu}{\sigma}=\frac{63-60}{\sigma}=\frac{3}{\sigma}"